OpenStudy (anonymous):

To skate at Roller Heaven, each person must pay a membership fee (m) and a fee for each session (f). Art attended 5 sessions and paid $25 in all. Later, when he had attended 10 sessions, he paid $40 in all. How much is the membership fee and how much are the sessions. Write the solution as an ordered pair.

OpenStudy (anonymous):

25 = m + 5f 40 = m +10f -40 = -m - 10f 25 = m + 5f 15 = 5f f = 3 25 = m +15 m = 10